Determine the retirement savings contributions credit in each of the following cases:

  1. A married couple filing jointly with modified AGI of $37,500 and an IRA contribution of $1,600.

  2. A married couple filing jointly with modified AGI of $58,000 and an IRA contribution of $1,500.

  3. A head of household taxpayer with modified AGI of $33,000 and Roth IRA contribution of $2,000.

  4. A single taxpayer with modified AGI of $12,000 and an IRA contribution of $2,300.
Answer #1

Solution:

1. Retirement savings cont. credit = $1,600×50% = $800

Since, married couple filing jointly will receive a credit of 50% of their contributions if their AGI is $38,500 or less

2. Retirement savings cont. credit = $1,500×10% = $150

Since, married couple filing jointly will receive a credit of 10% of their contributions if their AGI is between $41,501 to $64,000

3. Retirement savings cont. credit = $2,000×10% = $200

Head of household will receive 10% credit if their AGI is between $31,126 to $48,000

4. Retirement savings cont. credit = $2,300×50% = $1,150

Single filers will receive 50% credit if AGI is $19,250 or less

Note: If you have any query, ask me in the comment section. Please give thumbs(ratings) !!!